OpenStudy (anonymous):

A stone is thrown upward from ground level. The initial speed is 128 feet per second. How high will it go?

OpenStudy (anonymous):

use kinematic equation: \[v_f^2=v_i^2+2a(y_f-y_i)\] vf is zero at max hight, vi u know: -vi^2/2a=(yf-yi) yf-yi=-vi^2/2(-9.8) h=835.91 feet

OpenStudy (anonymous):

hold on leme give u the answer chocies

OpenStudy (anonymous):

a) 288 b) 272 c) 256 d) 248 e) 264

OpenStudy (anonymous):

non of above, it is 825.91 feet or 331 meters

OpenStudy (anonymous):

well taht is not an asnwer choice so it has to b one of teh follwoeing